A projectile is fired in such away that its horizontal range is equal to its maximum height what is the angle of projection ?

1
Expert's answer
2022-04-28T16:42:32-0400
"H_{max}=L_{max}"

"\frac{v_0^2\sin^2\theta}{2g}=\frac{v_0^2\sin2\theta}{g}"

"\tan\theta=4"

"\theta=76^\circ"